# Physics-inspired Generative Models in Medical Imaging: A Review

### v1

@@ 3.1.3 Nuclear Imaging
As another innovative approach for PET imaging, Chen et al. [17] introduced the 2.5D Multi-view Averaging Diffusion Model (MADM). This model converts non-attenuation-corrected low-dose PET (NAC-LDPET) into attenuation-corrected standard-dose PET (AC-SDPET), lowering the tracer dose and eliminating the need for CT-based attenuation correction. MADM uses separate diffusion models for axial, coronal, and sagittal views, averaging outputs to produce high-quality 3D images. The 2.5D architecture reduces computational demands and mitigates 2D discontinuities. In their experiments, MADM outperformed the competitive methods, offering high-quality 3D PET images with reduced radiation exposure and improved computational efficiency.
To further address the challenges of generalizability across different image noise levels, acquisition protocols, and patient populations in low-dose PET imaging, Xie et al. [105] developed DDPET-3D, a dose-aware diffusion model specifically for 3D low-dose PET imaging. Traditional diffusion models struggled with consistent 3D reconstructions and generalization across varying noise levels, often resulting in visually appealing but distorted images. DDPET-3D was extensively evaluated using 9,783 F-FDG studies from 1,596 patients across four medical centers with different scanners and clinical protocols. The model demonstrated strong generalizability to different low-dose levels, scanners, and protocols. Reader studies by nuclear medicine physicians confirmed that DDPET-3D produced superior denoised results comparable to or better than 100% full-count images and previous deep learning baselines. Real low-dose scans were included to showcase the clinical potential of DDPET-3D, indicating the feasibility of achieving high-quality low-dose PET images.
@@ 3.2 Medical Image Generation
Generating synthetic images through advanced generative models is a crucial area that addresses the scarcity of annotated medical datasets and ethical concerns associated with using real patient data. Techniques based on physics-informed models and their derivatives are particularly effective at producing high-quality, diverse synthetic images that accurately replicate human anatomy and disease variations, providing valuable resources for training algorithms and testing new technologies [49].
@@ 3.2.1 High-Quality Image Synthesis
Diffusion models have revolutionized medical image synthesis by addressing dataset scarcity, memory constraints, and computational complexity. These models generate high-quality images realistically, providing essential resources for developing and testing medical imaging algorithms [73].
To address the limited availability of 3D medical image datasets, Zhu et al. [122] developed GEM-3D, a generative approach using conditional diffusion models to synthesize and enhance 3D medical images. GEM-3D begins with a 2D slice containing patient-specific information and uses a 3D segmentation mask to guide the generation process. By decomposing 3D images into masks and patient prior information, GEM-3D enables versatile image generation, counterfactual synthesis, and dataset enhancement. This approach improves the quality and variability of generated 3D images, tackling high-quality 3D generation and condition decoupling for re-sampling data. The key innovation lies in decomposing segmentation masks and patient information for data augmentation, improving distribution coverage.
@@
To ease the tension between high-dimensional data and limited GPU memory [121],  [32] introduced wavelet diffusion models (WDM), a wavelet-based medical image synthesis framework that applies diffusion models to wavelet-decomposed images. WDM generates high-resolution 3D medical images by creating synthetic wavelet coefficients and then applying inverse discrete wavelet transform (IDWT). By decomposing an input image into a limited number of wavelet coefficients, concatenated into a single target matrix for the diffusion model, WDM maintains the standard architecture’s width through an initial convolution. Operating in the wavelet domain reduces spatial dimensions significantly, allowing for shallower networks, fewer computations, and reduced memory usage. During training, a corrupted sample is drawn, and the diffusion model approximates the original wavelet coefficients in terms of mean squared error. Inference starts with a sample from a normal distribution, and goes through reverse diffusion to reconstruct a final image. In the experiments, WDM outperformed the state-of-the-art methods in high-resolution 3D medical image generation.
In another study, Huang et al. [48] introduced a cascaded amortized latent diffusion model (CA-LDM) to synthesize high-resolution OCT volumes efficiently. This model employs non-holistic auto-encoders (NHAE) to map between low-resolution latent spaces and high-resolution volumes, achieving volumetric super-resolution slice-by-slice to manage the memory constraint. Cascaded diffusion processes further enhance this memory efficiency by first synthesizing the 3D global latent representation, followed by a 2D slice-wise refinement for high-resolution details. This approach effectively balances memory and speed in a task-specific way.
@@ 3.2.3 Data Imbalance and Image Robustness
Chambon et al. [13] introduced a vision-language foundation model called RoentGen to generate synthetic chest X-rays (CXR) based on text prompts. It leverages a pre-trained Latent Diffusion Model (LDM) fine-tuned on publicly available CXRs and radiology reports for the creation of diverse, high-quality images guided by radiology-specific text prompts. This method enhances image classification tasks through data augmentation. RoentGen’s architecture includes a variational autoencoder (VAE), a conditional denoising U-Net, and a CLIP text encoder, all fine-tuned with the MIMIC-CXR dataset. The model samples noise from a Gaussian distribution and predicts the original noise using the U-Net and encoded text prompts.
Another contribution to this field is Cheff, a cascaded latent diffusion model introduced by Weber et al. [97]. Cheff generates highly realistic 1-megapixel chest radiographs and addresses class imbalances by leveraging “MaCheX," the largest open collection of chest X-ray datasets, which supports report-to-chest-X-ray generation [88]. This methodology integrates auto-encoders and super-resolution diffusion models, setting a new standard in radiological image synthesis.
@@ 3.3 Medical Image Analysis
To shorten prediction time of the diffusion model, Zhang et al. [117] presented CTS, a consistency-based medical image segmentation model. By condensing sampling into a single iteration, CTS accelerates training and prediction while maintaining the generation quality [86]. This model uses multi-scale feature supervision and customized loss functions to improve segmentation and convergence. It integrates multi-scale signals into the UNet model and iteratively refines parameters, capitalizing contextual information.
Finally, FlowSDF [9] is another image-guided conditional flow matching framework through the signed distance function (SDF), facilitating an implicit distribution of segmentation masks. By leveraging SDF, FlowSDF achieves more natural distortions and high-quality sampling from the segmentation mask distribution of interest, enhancing predictive robustness with statistical tools like uncertainty maps. This method employs ODE-based flow matching to model probability flow, ensuring a smoother distribution and natural transitions between class modes. FlowSDF integrates image-guided flow matching with SDFs, enabling noise-injection for generating multiple segmentation maps and uncertainty quantification, thus improving model robustness and interpretability.
@@ 4 FUTURE DIRECTIONS
Physics-inspired generative models excel at closely matching the distribution of real images. This has not only paved the way for various medical imaging applications (as outlined in the previous section), but also offers potential for future developments along several research directions. Since hot topics for DDPMs are comprehensively covered in [54], this section focuses on other types of physics-inspired generative models, as well as more general aspects such as unification of generative models and performance considerations.
4.1 Unification of Score-matching Generative Models
DDPMs are grounded on the concept of non-equilibrium thermodynamics, specifically employing a diffusion-like process to progressively shape random noise into coherent data. PFGM are based on electrostatics, where the Poisson equation is used to describe the evolution of particles under the influence of electrostatic force. Accordingly, the question arises whether other physical phenomena can also give rise to generative models, or in other words, whether a duality between a physical phenomenon and a respective score-matching generative model exists [65].  Liu et al. [65] show that this can indeed be the case, under several restrictions on the physical processes. This opens up a wide range of new options that take advantage of physical processes for advancing medical imaging research.
4.2 Integration with Vision-Language Models
Whilst diffusion models have shown excellent performance on various image generation tasks, Large Language Models (LLMs) are dominant for generative tasks in natural language processing (NLP). Nevertheless, if a dedicated visual tokenizer that maps pixel-space inputs to discrete tokens can be effectively analyzed by LLMs, the resultant vision-language models may even outperform diffusion models on image and video generation benchmarks including ImageNet and Kinetics [111].
Instead of using either a diffusion model or a large large model, an increasing number of research efforts are capitalizing on both types of models synergistically. The canonical way of obtaining vision-language models (VLMs), or more generally multimodal models, would be to combine several models; for example, an image encoder to understand images, an LLM to generate text, and a diffusion model to generate images. A very interesting recent development is Chameleon [90], which represents a new approach of early-fusion token-based mixed-modal models designed to comprehend and produce images and text in any order. The flexibility of such models may prove useful in the field of medical imaging.
@@ 4.2 Integration with Vision-Language Models
Therefore, extensive research efforts are under way address the questions how to link both models and how to take advantage of the strengths of each model for generative tasks. In this context, there are several studies on different types of generative models. For instance, in [60], a diffusion model for text-to-image scene generation is enhanced with a large language model that generates a scene layout with captioned bounding boxes to handle more complex prompts. In [19] an autoencoder is enhanced with high-level features from a pre-trained LLM and its vocabulary. Regarding PFGM and PFGM++, their integration with large language models is largely untouched so far. Therefore, further research is clearly needed.
4.3 Novel Applications of Generative Models
Physics-inspired generative models have shown great potential in the context of various medical imaging modalities and associated tasks. It is expected that these models will be adapted to novel imaging modalities, such as phase-contrast and dark-field imaging techniques, spectral imaging, or multi-modal and multi-contrast approaches. Furthermore, generalist diffusion models are attractive, because these foundation models will be able to simultaneously cover multiple medical image domains and handle diverse clinical tasks. By taking advantage of multiple domains simultaneously, more robust performance on clinical tasks with scarce data is to be expected using the AI generalist approach.
Despite these potentials, there are several limitations regarding the score-matching generative models in medical imaging. Medical image data is always linked with protected health information (PHI), and protecting PHI is a critical aspect that needs to be taken care of. Although some of the proposed diffusion models could have a large clinical impact in medical imaging and image analysis, in many cases limitations arise due to the lack of data availability limited by the PHI concerns. Moreover, different imaging modalities, varying quality of scans and diverse patient demographics are inherent characteristics of medical image data. Furthermore, insufficient computing resources is an obstacle for diffusion model training when addressing non-trivial datasets, given an extensive computational cost of diffusion models in general. Memory efficient methods, such as low-rank adaptation (LoRA) [47], offer some democratization of the field by enabling users to adapt diffusion models to their domain of interest at a lower computational cost. Finally, general theoretical insights regarding score-matching generative models would be generally desirable and particularly in medical imaging areas. Further theoretical insights and practical solutions are anticipated to address potential biases, instability, hallucination, and sub-optimality in training quality and efficiency.
@@ 5 CONCLUSION
In this article, we have provided a general overview of different physics-inspired generative models and associated acceleration techniques. Furthermore, we have highlighted their applications in different medical imaging scenarios, including medical image reconstruction, image generation, and image analysis. Finally, we have discussed future research directions, such as unification of generative models, integration with language-vision models, new applications of generative models, including foundation models, as well as challenges to overcome. Clearly, there is immense potential for physics-inspired generative models and great opportunities to solve the current problems and redefine medical imaging.

### v2

@@ 3.1.3 Nuclear Imaging
As another innovative approach for PET imaging, Chen et al. [18] introduced the 2.5D Multi-view Averaging Diffusion Model (MADM). This model converts non-attenuation-corrected low-dose PET (NAC-LDPET) into attenuation-corrected standard-dose PET (AC-SDPET), lowering the tracer dose and eliminating the need for CT-based attenuation correction. MADM uses separate DMs for axial, coronal, and sagittal views, averaging outputs to produce high-quality 3D images. The 2.5D architecture reduces computational demands and mitigates 2D discontinuities. In their experiments, MADM outperformed the competitive methods, offering high-quality 3D PET images with reduced radiation exposure and improved computational efficiency.
To further address the challenges of generalizability across different image noise levels, acquisition protocols, and patient populations in low-dose PET imaging, Xie et al. [110] developed DDPET-3D, a dose-aware diffusion model specifically for 3D low-dose PET imaging. Traditional DMs struggled with consistent 3D reconstructions and generalization across varying noise levels, often resulting in visually appealing but distorted images. DDPET-3D was extensively evaluated using 9,783 F-FDG studies from 1,596 patients across four medical centers with different scanners and clinical protocols. The model demonstrated strong generalizability to different low-dose levels, scanners, and protocols. Reader studies by nuclear medicine physicians confirmed that DDPET-3D produced superior denoised results comparable to or better than 100% full-count images and previous deep learning baselines. Real low-dose scans were included to showcase the clinical potential of DDPET-3D, indicating the feasibility of achieving high-quality low-dose PET images.
@@ 3.2 Medical Image Generation
Generating synthetic images through advanced generative models is a crucial area that addresses the scarcity of annotated medical datasets and ethical concerns associated with using real patient data. Techniques based on physics-informed models and their derivatives, explored in Section 2, are particularly effective at producing high-quality, diverse synthetic images that accurately replicate human anatomy and disease variations, providing valuable resources for training algorithms and testing new technologies [51].
@@ 3.2.1 High-Quality Image Synthesis
DMs have revolutionized medical image synthesis by addressing dataset scarcity, memory constraints, and computational complexity. These models generate high-quality images realistically, providing essential resources for developing and testing medical imaging algorithms [78].
To address the limited availability of 3D medical image datasets, Zhu et al. [130] developed GEM-3D, a generative approach using conditional DMs to synthesize and enhance 3D medical images. GEM-3D begins with a 2D slice containing patient-specific information and uses a 3D segmentation mask to guide the generation process. By decomposing 3D images into masks and patient prior information, GEM-3D enables versatile image generation, counterfactual synthesis, and dataset enhancement. This approach improves the quality and variability of generated 3D images, tackling high-quality 3D generation and condition decoupling for re-sampling data. The key innovation lies in decomposing segmentation masks and patient information for data augmentation, improving distribution coverage.
@@
To ease the tension between high-dimensional data and limited GPU memory [129],  [33] introduced wavelet diffusion models (WDMs), a wavelet-based medical image synthesis framework that applies DMs to wavelet-decomposed images. WDM generates high-resolution 3D medical images by creating synthetic wavelet coefficients and then applying inverse discrete wavelet transform (IDWT). By decomposing an input image into a limited number of wavelet coefficients, concatenated into a single target matrix for the diffusion model, WDM maintains the standard architecture’s width through an initial convolution. Operating in the wavelet domain reduces spatial dimensions significantly, allowing for shallower networks, fewer computations, and reduced memory usage. During training, a corrupted sample is drawn, and the diffusion model approximates the original wavelet coefficients in terms of mean squared error. Inference starts with a sample from a normal distribution, and goes through reverse diffusion to reconstruct a final image. In the experiments, WDM outperformed the state-of-the-art methods in high-resolution 3D medical image generation.
In another study, Huang et al. [50] introduced a cascaded amortized latent diffusion model (CA-LDM) to synthesize high-resolution OCT volumes efficiently. This model employs non-holistic auto-encoders (NHAE) to map between low-resolution latent spaces, as mentioned in Section 2.6, and high-resolution volumes, achieving volumetric super-resolution slice-by-slice to manage the memory constraint. Cascaded diffusion processes further enhance this memory efficiency by first synthesizing the 3D global latent representation, followed by a 2D slice-wise refinement for high-resolution details. This approach effectively balances memory and speed in a task-specific way.
@@ 3.2.3 Data Imbalance and Image Robustness
Chambon et al. [14] introduced a vision-language foundation model called RoentGen to generate synthetic chest X-rays (CXR) based on text prompts. It leverages a pre-trained LDM fine-tuned on publicly available CXRs and radiology reports for the creation of diverse, high-quality images guided by radiology-specific text prompts. This method enhances image classification tasks through data augmentation. RoentGen’s architecture includes a VAE, a conditional denoising U-Net, and a CLIP text encoder. The U-Net and text encoder are fine-tuned on the MIMIC-CXR dataset while the VAE is kept frozen. The model samples noise from a Gaussian distribution and predicts the original noise using the U-Net and encoded text prompts.
Another contribution to this field is Cheff, a cascaded latent diffusion model introduced by Weber et al. [102]. Cheff generates highly realistic 1-megapixel chest radiographs and addresses class imbalances by leveraging “MaCheX", the largest open collection of chest X-ray datasets, which supports report-to-chest-X-ray generation [93]. This methodology integrates auto-encoders and super-resolution DMs, setting a new standard in radiological image synthesis.
@@ 3.3 Medical Image Analysis
To shorten prediction time of the diffusion model, Zhang et al. [122] presented CTS, a consistency-based medical image segmentation model, as noted in Section 2.7. By condensing sampling into a single iteration, CTS accelerates training and prediction while maintaining the generation quality [91]. This model uses multi-scale feature supervision and customized loss functions to improve segmentation and convergence. It integrates multi-scale signals into the UNet model and iteratively refines parameters, capitalizing contextual information.
In response to the limitations of 2D methods, which often miss crucial out-of-slice information, Kim and Park [58]. introduced a refined LDM with a novel switchable block, the multiple switchable spatially adaptive normalization (MS-SPADE). This innovation enables effective 3D medical image translation without patch cropping, allowing for the translation of multiple target modalities from a single source. The MS-SPADE block dynamically adjusts source latents to match target styles, improving fidelity and reducing the need for multiple models, thereby addressing practical challenges like high scan costs and limited availability in clinical settings.
Furthermore, Arslan et al. [8] address key challenges in DDMs used for medical image translation, particularly the weak guidance from source modalities and the poor alignment between denoising and the necessary transformations from source to target that reduce effectiveness [41]. They present the Self-consistent Recursive Diffusion Bridge (SelfRDB). This model markedly betters traditional DDMs by introducing a forward process with a strategy that gradually increases noise levels, improving generalization and boosting the transfer of information between modalities. SelfRDB stands out from typical DMs by using a special recursive estimation method in its backward steps, which permits ongoing improvement of the target image while keeping the original tissue details intact. Extensive tests on multi-contrast MRI and MRI-CT translations show that SelfRDB surpasses current GAN [128] and DMs and is the first to employ a soft-prior on the source modality and a consistent recursive estimation to refine the accuracy of backward diffusion steps.
Zheng et al. [127] presented the Deformation-Recovery Diffusion Model (DRDM), a novel approach that shifts from traditional intensity-based synthesis [84] to deformation-based methods [123], using a mechanism that generates topologically-preserving multi-scale Deformation Velocity Fields (DVFs). This model enhances the interpretability and anatomical accuracy of synthetic medical images by creating physically plausible deformation fields. Capable of managing instance-specific deformations without reference images, the DRDM supports applications such as data augmentation for few-shot learning and synthetic training for image registration. This advancement enriches synthetic medical image generation and aids critical downstream tasks without the need for external annotations.
Finally, FlowSDF [10] is another image-guided conditional flow matching framework through the signed distance function (SDF), facilitating an implicit distribution of segmentation masks. By leveraging SDF, FlowSDF achieves more natural distortions and high-quality sampling from the segmentation mask distribution of interest, enhancing predictive robustness with statistical tools like uncertainty maps. This method employs ODE-based flow matching [65, 68] to model probability flow, ensuring a smoother distribution and natural transitions between class modes. FlowSDF integrates image-guided flow matching with SDFs, enabling noise-injection for generating multiple segmentation maps and uncertainty quantification, thus improving model robustness and interpretability.
@@ 4 FUTURE DIRECTIONS
Physics-inspired GMs excel at closely matching the distribution of real images. This has not only paved the way for various medical imaging applications (as outlined in the previous section), but also offers potential for future developments along several research directions. Since hot topics for DDPMs are comprehensively covered in [56], this section focuses on other types of physics-inspired GMs, as well as more general aspects such as unification of GMs and performance considerations.
4.1 Unification of SDMs
DDPMs (cf. Section 2.1) are grounded on the concept of non-equilibrium thermodynamics, specifically employing a diffusion-like process to progressively shape random noise into coherent data. PFGMs (cf. Section 2.4) are based on electrostatics, where the Poisson equation is used to describe the evolution of particles under the influence of electrostatic force. Accordingly, the question arises whether other physical phenomena can also give rise to GMs, or in other words, whether a duality between a physical phenomenon and a respective “score-matching” GM exists [70].  Liu et al. [70] show that this can indeed be the case, under several restrictions on the physical processes. This opens up a wide range of new options that take advantage of physical processes for advancing medical imaging research.
4.2 Integration with VLMs
Whilst DMs have shown excellent performance on various image generation tasks(both in medical imaging as detailed in Section 3, as well as various other fields), Large Language Models (LLMs) are dominant for generative tasks in natural language processing (NLP). Nevertheless, if a dedicated visual tokenizer that maps pixel-space inputs to discrete tokens can be effectively analyzed by LLMs, the resultant vision-language models may even outperform DMs on image and video generation benchmarks including ImageNet and Kinetics [116].
Instead of using either a DM or a LLM, an increasing number of research efforts are capitalizing on both types of models synergistically. The canonical way of obtaining VLMs, or more generally multimodal models, would be to combine several models; for example, an image encoder to understand images, an LLM to generate text, and a DM to generate images. A very interesting recent development is Chameleon [95], which represents a new approach of early-fusion token-based mixed-modal models designed to comprehend and produce images and text in any order. The flexibility of such models may prove useful in the field of medical imaging.
@@ 4.2 Integration with VLMs
Therefore, extensive research efforts are under way to explore how to link both models and how to take advantage of the strengths of each model for generative tasks. In this context, there are several studies on different types of GMs. For instance, in [63], a DM for text-to-image scene generation is enhanced with a LLM that generates a scene layout with captioned bounding boxes to handle more complex prompts. In [20] an autoencoder is enhanced with high-level features from a pre-trained LLM and its vocabulary. Regarding PFGMs and PFGM++ (cf. Sections 2.4 and 2.5), their integration with LLMs is largely untouched so far. Therefore, further research is clearly needed.
4.3 Novel Applications of GMs
Physics-inspired GMs have shown great potential in the context of various medical imaging modalities (cf. Section 3.1) and associated tasks (cf. Sections 3.2 and 3.3). It is expected that these models will be adapted to novel imaging modalities, such as phase-contrast and dark-field imaging techniques, spectral imaging, or multi-modal and multi-contrast approaches. Furthermore, generalist DMs are attractive, because these foundation models will be able to simultaneously cover multiple medical image domains and handle diverse clinical tasks. By taking advantage of multiple domains simultaneously, more robust performance on clinical tasks with scarce data is to be expected using the AI generalist approach.
Despite these potentials, there are several limitations regarding the score-matching generative models in medical imaging. Medical image data is always linked with protected health information (PHI), and protecting PHI is a critical aspect that needs to be taken care of. Although some of the proposed DMs could have a large clinical impact in medical imaging and image analysis, in many cases limitations arise due to the lack of data availability limited by the PHI concerns. This is particularly evident in applications such as personalized medicine, rare diseases, longitudinal studies, and multi-institutional collaborations. Moreover, different imaging modalities, varying quality of scans and diverse patient demographics are inherent characteristics of medical image data. Furthermore, insufficient computing resources is an obstacle for DM training when addressing non-trivial datasets, given an extensive computational cost of DMs in general. Memory efficient methods, such as Low-Rank Adaptation (LoRA) [49], offer some democratization of the field by enabling users to adapt DMs to their domain of interest at a lower computational cost. Finally, general theoretical insights regarding score-matching GMs would be generally desirable and particularly in medical imaging areas. Further theoretical insights and practical solutions are anticipated to address potential biases, instability, hallucination, and sub-optimality in training quality and efficiency.
@@ 5 CONCLUSION
In this article, we have provided a general overview of different physics-inspired GMs and associated acceleration techniques. Furthermore, we have highlighted their applications in different medical imaging scenarios, including medical image reconstruction, image generation, and image analysis. Finally, we have discussed future research directions, such as unification of GMs, integration with VLMs, new applications of GMs, including foundation models, as well as challenges to overcome. Clearly, there is immense potential for physics-inspired GMs and great opportunities to solve the current problems and redefine medical imaging.